# Using $\Lambda_{b}^{0}(6146)$ and $\Lambda_{b}^{0}(6152)$ as probes to investigate possible $\bar{B}^{*}N$ and $\bar{D}^{*}N$ molecules

### v1

@@ Abstract
Heavy quark symmetry can help us identify the internal structure of hadrons and predict new particles. In this study, we examine the strong decay modes of the observed $\Lambda_{b}^{0}(6146)$ and $\Lambda_{b}^{0}(6152)$, assuming these two states are molecular states primarily composed of $\bar{B}^{*}N$ component. The partial decay widths of the $\bar{B}^{*}N$ molecular state into the $\pi\Sigma_{b}$ and $\pi\Sigma_{b}^{*}$ final states through hadronic loops are calculated using effective Lagrangians. Our results, when compared with LHCb observations, support the interpretation of $\Lambda_{b}^{0}(6146)$ as a molecule primarily composed of $\bar{B}^{*}N$ components. However, the decay width of $\Lambda_{b}^{0}(6152)$ cannot be accurately reproduced within the molecular state framework. Based on the above results and heavy quark symmetry, we predict the existence of $\bar{B}^{*}N$ molecular states with $J^{p}=5/2^{+}$, which are the heavy quark spin symmetry partners of $\Lambda_{b}(6146)$, with masses in the range of 6195-6200 MeV. And the main decay is $\pi\Sigma_{b}^{*}$ channel. Moreover, there must existence of a $\bar{D}^{*}N$ molecule with $J^{p}=3/2^{+}$, possible corresponding to the experimentally observed $\Lambda_{c}(2860)^{+}$. If $\Lambda_{c}(2880)^{+}$ is indeed the heavy quark flavor symmetry partner of $\Lambda_{b}(6152)$, it would exhibit a conventional three-quark structure. Therefore, we also propose the search for a $\bar{D}^{*}N$ molecule with a spin-parity of $J^{p}=5/2^{+}$, which would be the heavy-quark spin partner state of $\Lambda_{c}(2860)^{+}$. It should be noted that these baryons may be mixed states, containing both molecular and three-quark components. These results can aid experiments in exploring the internal structure of these baryons.
@@ 1 Introduction
$\Lambda_{\_}c(2595)^{+}$ $1/2^{-}$ $\bar{D}N~{}$Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina ; Tolos:2004yg ; Hofmann:2005sw ; Mizutani:2006vq ; Jimenez-Tejero:2009cyn $\Lambda_{\_}b(5912)^{0}$ $1/2^{-}$ $\bar{B}N$Kawakami:2018olq ; Lu:2014ina ; Garcia-Recio:2012lts ; Liang:2014eba ; Huang:2021ave
$\Lambda_{\_}c(2625)^{+}$ $3/2^{-}$ $\bar{D}^{*}N~{}$Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina $\Lambda_{\_}b(5920)^{0}$ $3/2^{-}$ $\bar{B}^{*}N$Kawakami:2018olq ; Lu:2014ina ; Garcia-Recio:2012lts ; Liang:2014eba ; Huang:2021ave
$\Lambda_{\_}c(2860)^{+}$ $3/2^{+}$ $\bar{D}^{(*)}N~{}$? $\Lambda_{\_}b(6146)^{0}$ $3/2^{+}$ $\bar{B}^{(*)}N$ Jian:2022rln
$\Lambda_{\_}c(2880)^{+}$ $5/2^{+}$ $\bar{D}^{(*)}N~{}$? $\Lambda_{\_}b(6152)^{0}$ $5/2^{+}$ $\bar{B}^{(*)}N$ Jian:2022rln
A more significant reason is based on heavy quark symmetry Isgur:1991wq , which incorporates both heavy quark spin symmetry and heavy quark flavor symmetry. Experimentally, several charm-containing baryons $\Lambda_{\_}c^{+}$ have been observed ParticleDataGroup:2022pth . Among them, $\Lambda_{\_}c(2595)^{+}$ and $\Lambda_{\_}c(2625)^{+}$ can be interpreted as molecular states mainly composed of $\bar{D}N$ and $\bar{D}^{*}N$ components Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina ; Tolos:2004yg ; Hofmann:2005sw ; Mizutani:2006vq ; Jimenez-Tejero:2009cyn , respectively (see Tab. 1 ), forming heavy quark spin symmetry (HQSS) doublet states. Simultaneously, heavy quark spin symmetry also implies the existence of other high spin-parity HQSS doublet states corresponding to the $\bar{D}^{(*)}N$ system, with their spin-parity being $J^{p}=3/2^{+}$ and $J^{p}=5/2^{+}$, respectively Yamaguchi:2014era . After carefully examining the PDG ParticleDataGroup:2022pth , it is found that these two molecular states may correspond to $\Lambda_{\_}c(2860)^{+}$ and $\Lambda_{\_}c(2880)^{+}$, respectively. However, there is currently no research on whether they can be considered hadrons containing molecular components. According to heavy quark flavor symmetry, we propose a scheme to investigate whether they can be considered as $\bar{D}^{(*)}N$ molecules, which involves studying their heavy quark flavor partners. Thus, we hope that $\bar{B}^{(*)}N$ molecules exist.
@@ 3 Results and discussions
According to the heavy quark flavor symmetry, the observed $\Lambda_{\_}b(6146)$ can be considered a $p$-wave $\bar{B}N$ molecular state, suggesting the existence of a $\bar{D}^{*}N$ molecule with $J^{p}=3/2^{+}$. This $\bar{D}^{*}N$ molecule could correspond to the experimentally observed $\Lambda_{\_}c(2860)^{+}$. If $\Lambda_{\_}c(2880)^{+}$ is indeed the heavy quark flavor symmetry partner of $\Lambda_{\_}b(6152)$, it would possess a conventional three-quark structure. Consequently, we suggest seeking a $\bar{D}^{*}N$ molecule with a spin-parity of $J^{p}=5/2^{+}$, which is the heavy-quark spin partner state of $\Lambda_{\_}c(2860)^{+}$.
@@ 4 Summary
Under the guidance of heavy quark symmetry, $\Lambda_{\_}b(6146)$ has the properties of a $P$-wave $N\bar{B}^{*}$ molecule, suggesting the existence of a $\bar{D}^{*}N$ molecule with spin $J^{P}=3/2^{+}$, which resonates with the experimentally observed $\Lambda_{\_}c(2860)^{+}$. If $\Lambda_{\_}c(2880)^{+}$ is confirmed to be the heavy quark flavor symmetry partner of $\Lambda_{\_}b(6152)$, it is expected to have a typical three-quark structure. In the mass range of 6195$\sim$6200 MeV, we believe that there may be an $\bar{B}^{*}N$ molecular state with $J^{P}=5/2^{+}$ as the heavy quark spin symmetry partner of $\Lambda_{\_}b(6146)$. Therefore, we suggest conducting experiments to search for this $b$-quark-containing baryon and its heavy quark flavor symmetry partner. This also tell us that not all pairs of hadrons containing a heavy quark (c or b), with a mass difference of only a few MeV and almost equal decay widths, necessarily form heavy quark spin symmetry doublet states.

### v2

# Using $\Lambda_{b}^{0}(6146)$ and $\Lambda_{b}^{0}(6152)$ as probes to investigate possible $\bar{B}^{*}N$ and $D^{*}N$ molecules

@@ Abstract
Heavy quark symmetry can help us identify the internal structure of hadrons and predict new particles. In this study, we examine the strong decay modes of the observed $\Lambda_{b}^{0}(6146)$ and $\Lambda_{b}^{0}(6152)$, assuming these two states are molecular states primarily composed of $\bar{B}^{*}N$ component. The partial decay widths of the $\bar{B}^{*}N$ molecular state into the $\pi\Sigma_{b}$ and $\pi\Sigma_{b}^{*}$ final states through hadronic loops are calculated using effective Lagrangians. Our results, when compared with LHCb observations, support the interpretation of $\Lambda_{b}^{0}(6146)$ as a molecule primarily composed of $\bar{B}^{*}N$ components. However, the decay width of $\Lambda_{b}^{0}(6152)$ cannot be accurately reproduced within the molecular state framework. Based on the above results and heavy quark symmetry, we predict the existence of $\bar{B}^{*}N$ molecular states with $J^{p}=5/2^{+}$, which are the heavy quark spin symmetry partners of $\Lambda_{b}(6146)$, with masses in the range of 6195-6200 MeV. And the main decay is $\pi\Sigma_{b}^{*}$ channel. Moreover, there must existence of a $D^{*}N$ molecule with $J^{p}=3/2^{+}$, possible corresponding to the experimentally observed $\Lambda_{c}(2860)^{+}$. If $\Lambda_{c}(2880)^{+}$ is indeed the heavy quark flavor symmetry partner of $\Lambda_{b}(6152)$, it would exhibit a conventional three-quark structure. Therefore, we also propose the search for a $D^{*}N$ molecule with a spin-parity of $J^{p}=5/2^{+}$, which would be the heavy-quark spin partner state of $\Lambda_{c}(2860)^{+}$. It should be noted that these baryons may be mixed states, containing both molecular and three-quark components. These results can aid experiments in exploring the internal structure of these baryons.
@@ 1 Introduction
$\Lambda_{\_}c(2595)^{+}$ $1/2^{-}$ $DN~{}$Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina ; Tolos:2004yg ; Hofmann:2005sw ; Mizutani:2006vq ; Jimenez-Tejero:2009cyn $\Lambda_{\_}b(5912)^{0}$ $1/2^{-}$ $\bar{B}N$Kawakami:2018olq ; Lu:2014ina ; Garcia-Recio:2012lts ; Liang:2014eba ; Huang:2021ave
$\Lambda_{\_}c(2625)^{+}$ $3/2^{-}$ $D^{*}N~{}$Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina $\Lambda_{\_}b(5920)^{0}$ $3/2^{-}$ $\bar{B}^{*}N$Kawakami:2018olq ; Lu:2014ina ; Garcia-Recio:2012lts ; Liang:2014eba ; Huang:2021ave
$\Lambda_{\_}c(2860)^{+}$ $3/2^{+}$ $D^{(*)}N~{}$? $\Lambda_{\_}b(6146)^{0}$ $3/2^{+}$ $\bar{B}^{(*)}N$ Jian:2022rln
$\Lambda_{\_}c(2880)^{+}$ $5/2^{+}$ $D^{(*)}N~{}$? $\Lambda_{\_}b(6152)^{0}$ $5/2^{+}$ $\bar{B}^{(*)}N$ Jian:2022rln
A more significant reason is based on heavy quark symmetry Isgur:1991wq , which incorporates both heavy quark spin symmetry and heavy quark flavor symmetry. Experimentally, several charm-containing baryons $\Lambda_{\_}c^{+}$ have been observed ParticleDataGroup:2022pth . Among them, $\Lambda_{\_}c(2595)^{+}$ and $\Lambda_{\_}c(2625)^{+}$ can be interpreted as molecular states mainly composed of $\bar{D}N$ and $\bar{D}^{*}N$ components Nieves:2019nol ; Garcia-Recio:2008rjt ; Nieves:2019kdh ; Guo:2019ytq ; Kawakami:2018olq ; Lu:2014ina ; Tolos:2004yg ; Hofmann:2005sw ; Mizutani:2006vq ; Jimenez-Tejero:2009cyn , respectively (see Tab. 1 ), forming heavy quark spin symmetry (HQSS) doublet states. Simultaneously, heavy quark spin symmetry also implies the existence of other high spin-parity HQSS doublet states corresponding to the $D^{(*)}N$ system, with their spin-parity being $J^{p}=3/2^{+}$ and $J^{p}=5/2^{+}$, respectively Yamaguchi:2014era . After carefully examining the PDG ParticleDataGroup:2022pth , it is found that these two molecular states may correspond to $\Lambda_{\_}c(2860)^{+}$ and $\Lambda_{\_}c(2880)^{+}$, respectively. However, there is currently no research on whether they can be considered hadrons containing molecular components. According to heavy quark flavor symmetry, we propose a scheme to investigate whether they can be considered as $D^{(*)}N$ molecules, which involves studying their heavy quark flavor partners. Thus, we hope that $\bar{B}^{(*)}N$ molecules exist.
@@ 3 Results and discussions
According to the heavy quark flavor symmetry, the observed $\Lambda_{\_}b(6146)$ can be considered a $p$-wave $\bar{B}^{*}N$ molecular state, suggesting the existence of a $D^{*}N$ molecule with $J^{p}=3/2^{+}$. This $D^{*}N$ molecule could correspond to the experimentally observed $\Lambda_{\_}c(2860)^{+}$. If $\Lambda_{\_}c(2880)^{+}$ is indeed the heavy quark flavor symmetry partner of $\Lambda_{\_}b(6152)$, it would possess a conventional three-quark structure. Consequently, we suggest seeking a $D^{*}N$ molecule with a spin-parity of $J^{p}=5/2^{+}$, which is the heavy-quark spin partner state of $\Lambda_{\_}c(2860)^{+}$.
@@ 4 Summary
Under the guidance of heavy quark symmetry, $\Lambda_{\_}b(6146)$ has the properties of a $P$-wave $N\bar{B}^{*}$ molecule, suggesting the existence of a $D^{*}N$ molecule with spin $J^{P}=3/2^{+}$, which resonates with the experimentally observed $\Lambda_{\_}c(2860)^{+}$. If $\Lambda_{\_}c(2880)^{+}$ is confirmed to be the heavy quark flavor symmetry partner of $\Lambda_{\_}b(6152)$, it is expected to have a typical three-quark structure. In the mass range of 6195$\sim$6200 MeV, we believe that there may be an $\bar{B}^{*}N$ molecular state with $J^{P}=5/2^{+}$ as the heavy quark spin symmetry partner of $\Lambda_{\_}b(6146)$. Therefore, we suggest conducting experiments to search for this $b$-quark-containing baryon and its heavy quark flavor symmetry partner. This also tell us that not all pairs of hadrons containing a heavy quark (c or b), with a mass difference of only a few MeV and almost equal decay widths, necessarily form heavy quark spin symmetry doublet states.